# Zero-Shot Statistical Tests for LLM-Generated Text Detection using Finite Sample Concentration Inequalities

### v3

@@ 3 Concentration bounds
In this section, we present our results in two parts. In Section 3.1, we provide concentration bounds to show that if the string is generated by model $A$, $\frac{1}{N}\sum_{n=1}^{N}Z_{n}$ converges to the average entropy of the string under $A$ with a high probability. In Section 3.2, we provide concentration bounds to show that if the string is generated by another model $B$, then $\frac{1}{N}\sum_{n=1}^{N}Z_{n}$ converges to the average cross-entropy of the string under $B$ and $A$ with a high probability. These concentration bounds are the backbones of the statistical tests that we design in section 4.
3.1 Same generative and evaluator models
Consider a string $\mathbf{Y}$ generated by model $A$ and we evaluate the text using the same model $A$. First recall that given a string $\mathbf{Y}_{N}$ and an evaluator model $A$, we define the random variable $Z_{n}=-\log\big{(}p^{A}_{n}(Y_{n})\big{)}$. Then,
$\mathbb{E}_{p_{n}^{A}}[Z_{n}]=-\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})\log p^{A}_{n}(y_{n}).$
Also, given the string $\mathbf{Y}_{N}$, since we have complete knowledge on the probabilities $p^{A}_{n}(y_{n})$, we have complete knowledge on the entropy $-\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})\log p^{A}_{n}(y_{n})=\mathbb{E}_{p_{n}^{A}}[Z_{n}]$. Then, we can find the following upper bound for $\mathbb{E}_{p_{n}^{A}}[Z_{n}]$.
Remark 1.
$\mathbb{E}_{p_{n}^{A}}[Z_{n}]=-\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})\log p^{A}_{n}(y_{n})\leq\log|\mathcal{X}|=\log(K).$
@@ 3.1 Same generative and evaluator models
By concavity of $-p^{A}_{n}(y_{n})\log p^{A}_{n}(y_{n})$, the value of its maximizer is $p^{*A}_{n}(y_{n})=\frac{1}{|\chi|},\quad\forall y_{n}\in\chi$. Thus,
$\mathbb{E}_{p_{n}^{A}}[Z_{n}]=-\sum_{y_{n}\in\chi}p_{n}^{A}(y_{n})\log p_{n}^{A}(y_{n})$$\leq$ $\sum_{y_{n}\in\chi}p_{n}^{*A}(y_{n})\log p_{n}^{*A}(y_{n})$
$=$ $|\chi|.\frac{1}{|\chi|}.\log|\chi|=\log(K).$
With $\mathbb{E}_{p_{n}^{A}}[Z_{n}]<\infty$, we define the zero-mean random variable $X_{n}=Z_{n}-\mathbb{E}_{p_{n}^{A}}[Z_{n}]$.
Lemma 1.
The random variable $S_{N}=\sum_{i=1}^{N}X_{i}$ forms a martingale.
@@
$Z_{n}$ and as a result $\mathbb{E}_{p_{n}^{A}}(y_{n})$ are positive random variables, and we showed that $\mathbb{E}_{p_{n}^{A}}[Z_{n}]<\log(K)$. Therefore, we have $\mathbb{E}\big{[}|X_{n}|\big{]}\leq\log(K)$. For a random variable to form a martingale, the following two properties need to be satisfied: (i) $\mathbb{E}\big{[}|S_{N+1}-S_{N}|\big{]}<\infty$, and (ii) $\mathbb{E}[S_{N+1}|S_{N}]=S_{N}$.
(i) is satisfied because $\mathbb{E}\big{[}|S_{N+1}-S_{N}|\big{]}=\mathbb{E}\big{[}|X_{N+1}|\big{]}\leq\log(K)<\infty$.
(ii) is satisfied because the martingale increments $X_{n}$ are, by definition, a zero-mean random variable conditioned on past tokens. ∎
Finally, we apply concentration bounds for martingales to provide finite sample guarantee for the convergence of the random variable $S_{N}/N$ to zero. A challenge in applying the common concentration bounds for martingales is that martingale increments are not necessarily bounded. We overcome this issue by showing that the martingale differences, while not bounded, admit a light tail. In particular, we show that the martingale differences are sub-exponential.
Definition 1.
If $\|X\|_{\psi_{1}}$ is finite, we say that $X$ is sub-exponential.
Lemma 2.
The sub-exponential norm for random variable $X_{n}$ equals $2\log(K)$.
@@
See Appendix A.1. ∎
As the last step, we apply the concentration bounds for martingales with sub-exponential increments to obtain the following concentration bound.
Theorem 1.
There exists a constant $c_{1}>0$ independent of the evaluator model $A$ such that for any $t>0$ we have
$\mathbb{P}\bigg{(}\frac{1}{N}|\sum_{n=1}^{N}X_{n}|\geq t\bigg{)}\leq 2\exp\bigg{[}-\frac{Nt}{c_{1}\log(K)}\min\bigg{(}1,\frac{t}{c_{1}\log(K)}\bigg{)}\bigg{]}$
Proof.
See Section A.2. ∎
Interpretation. Theorem 1 states that if a given string is generated by a model same as the evaluator (here, model $A$), then the log-perplexity of the string under $A$ converges to the average entropy of the string under $A$, except with an exponentially small probability in string length.
3.2 Different generative and evaluator models
In this section, we consider a string $\bf{Y}$ generated by model $B$ and want to evaluate our statistical test based on model $A$. To design the statistical test, first recall that given a string $\bf{Y}$ and an evaluator model $A$, we define the random variable $Z_{n}=-\log\big{(}p^{A}_{n}(y_{n})\big{)}$. Note that
$\mathbb{E}_{p^{B}_{n}}[Z_{n}]=\sum_{y_{n}\in\mathcal{X}}-p^{B}_{n}(y_{n})\log(p^{A}_{n}(y_{n}))=H(p^{B}_{n},p^{A}_{n}),$ (4)
where $H(p^{B}_{n},p^{A}_{n})$ is the cross entropy between the two distributions $p^{B}_{n}(.)$ and $p^{A}_{n}(.)$.
Note that, unlike the case analyzed in section 3.1, here, $\mathbb{E}_{p_{n}^{B}}[Z_{n}]$ is not necessarily finite.
For $\mathbb{E}_{p^{B}_{n}}[Z_{n}]$ to be infinite, as we can infer from equation 4, we must have that $p^{A}_{n}(y_{n})=0$ and $p^{B}_{n}(y_{n})>0$ for some $y_{n}\in\chi$. In this case, if the string includes such $y_{n}$, then we realize that the string is not generated by model $A$ with the probability of 1. This is a trivial case.
Yet, if the string does not include any such $y_{n}$, then we can update the probability distributions as
$\tilde{p}^{B}_{n}=\frac{p^{B}_{n}(y_{n})}{\sum_{y_{k}:p^{A}_{n}(y_{k})>0}p^{B}_{n}(y_{k})}.$
Without loss of generality, we can exclude the trivial case and assume that if $p^{B}_{n}(y_{n})>0$, then $p^{A}_{n}(y_{n})>0$, which yields that $\mathbb{E}_{p^{B}_{n}}[Z_{n}]=H(p^{B}_{n},p^{A}_{n})$ is finite. With $\mathbb{E}_{p_{n}^{B}}[Z_{n}]<\infty$, we define the zero-mean random variable $X_{n}=Z_{n}-\mathbb{E}_{p^{B}_{n}}[Z_{n}]$. For tractability, we make a parametric assumption on the probability laws $p^{A}_{k}(.)$ and $p^{B}_{k}(.)$.
Assumption 1.
For our two model analysis, we assume that there exists $\epsilon>0$ such that $p^{A}_{n}(y_{k}),p^{B}_{n}(y_{k})\notin(0,\epsilon)$.
Assumption 1 implies that models $A$ and $B$ either do not associate any probability to a token $y\in\mathcal{X}$, or they assign a probability of at least $\epsilon$. Our theoretical results depend only on $\log(\epsilon)$. Hence, our theoretical bounds primarily rely on a constant shift in the logarithmic scale.
It is noteworthy that Assumption 1 is not restrictive and is aligned with practice, as computers only allow for a limited range of representable numbers due to finite precision in floating-point arithmetic. Very small probabilities are either rounded to zero or set to a minimum threshold to maintain numerical stability in computations (Goldberg, 1991).
The first outcome of assumption 1 is that
$\mathbb{E}_{p^{B}_{n}}\big{[}|Z_{n}|\big{]}=\mathbb{E}_{p^{B}_{n}}\big{[}Z_{n}\big{]}=\sum_{y_{n}\in\mathcal{X}}-p^{B}_{n}(y_{n})\log(p^{A}_{n}(y_{n}))\leq-\log(\epsilon)$ (5)
Hence, with the same argument as in section 3.1, the random variable $S_{N}=\sum_{i=1}^{N}X_{i}$ forms a martingale. To apply the martingale concentration bounds, similar to the previous section, we first find the sub-exponential norm for the random variable $X_{n}$.
Lemma 3.
Proof.
See Section A.3. ∎
Theorem 2.
There exists a constant $c_{3}>0$ independent of models $A$ and $B$ such that for any $t>0$ we have
$\mathbb{P}\bigg{(}\frac{1}{N}|\sum_{n=1}^{N}X_{n}|\geq t\bigg{)}\leq 2\exp\bigg{[}-\frac{Nt}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{t}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}$
Proof.
See Section A.4. ∎
Interpretation. Theorem 2 states that if model $B$ generates the text, then, except with an exponentially small probability in string length, the log-perplexity of the string under model $A$ converges to the average cross-entropy of the string under $B$ and $A$.
4 Statistical test
Now we design our statistical tests using the results in Theorems 1 and 2 and then evaluate type I (false positive) and type II (false negative) errors. In particular, we consider a finite-length text with length $N$ generated by a model $M$. We first design tests for detection between different LLMs. In that, we have white-box access to all models in the hypothesis test. In Section 4.1, we design a simple statistical test that determines whether a text is generated by a model $A$ or another model $B$. Then, in Section 4.2, we extend our results to composite tests that determine whether a text is generated by a model that belongs to a set of models $\mathcal{A}$ or a model that belongs to a disjoint set of models $\mathcal{B}$. Finally, in Section 4.3, we study the case where we don’t have white-box access to all models in the hypothesis set (for example, a human produced the text) and design a composite statistical test to identify whether the text is generated by a model $A$ or not.
4.1 Simple statistical test for detection between two LLMs
Statistical test. Given a string $\mathbf{Y}_{N}$ with length $N$, we design a statistical test to detect whether model $A$ or model $B$ generated the text. The null hypothesis $\mathbf{H}_{0}$ is that the text $\mathbf{Y}_{N}$ is generated by $B$, and the alternative hypothesis $\mathbf{H}_{1}$ is that $\mathbf{Y}_{N}$ is generated by $A$. We first calculate the random variables $Z^{A}_{n}=:-\log\big{(}p^{A}_{n}(Y_{n})\big{)}$ and $Z^{B}_{n}=:-\log\big{(}p^{B}_{n}(Y_{n})\big{)}$, and then we calculate the sums $\frac{1}{N}\sum_{n=1}^{N}Z^{A}_{n}$, and $\frac{1}{N}\sum_{n=1}^{N}Z^{B}_{n}$. Our test rejects the null hypothesis $\mathbf{H}_{0}$ if
$\frac{1}{N}\sum_{n=1}^{N}Z^{A}_{n}<\frac{1}{N}\sum_{n=1}^{N}Z^{B}_{n}.$
Otherwise, our test accepts the null hypothesis.
Type I and type II errors. Type I error occurs when the test incorrectly concludes that the text is generated by the model $A$ when it is written by $B$, and Type II error happens when the test fails to identify that text is generated by the model $A$ and incorrectly concludes that it is generated by $B$.
To quantify our model’s type I and type II errors, we need to make the following (mild) assumption.
Assumption 2.
(minimum difference). We assume that if the generative and evaluator models are different, for an arbitrarily small positive $\epsilon_{1}>0$, we have
$\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\geq\epsilon_{1}.$
Assumption 2 ensures that the two models satisfy a minimum distance in terms of their KL divergence over the generated text. Note that KL divergence, by definition, is a non-negative value that demonstrates the distance between the two distributions over the next word for the two models. Our results show that the type I and type II errors of our statistical test are approximately $\exp\big{(}O(-N\epsilon_{1})\big{)}$, which indicates that even for small values of $\epsilon_{1}$ that can converge to zero with the length of text (for example, $\epsilon_{1}=O(N^{-1/2})$), our statistical test provides exponentially small type I and type II errors in the length of the text. Hence, our theoretical bounds only require that the two models do not impose the same probability distribution on the string.
Proposition 1.
If Assumptions 1 and 2 hold, then the type I and type II errors for our statistical test are upper bounded by
$2\exp\bigg{[}-\frac{N(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}+2\exp\bigg{[}-\frac{N\epsilon_{1}/2}{c_{1}\log(K)}\min\bigg{(}1,\frac{\epsilon_{1}/2}{c_{1}\log(K)}\bigg{)}\bigg{]}.$
with constants $c_{1}$, $c_{3}$, and $\epsilon$ as introduced in Theorems 1 and 2.
Proof.
See Section A.5. ∎
@@ 4.1 Simple statistical test for detection between two LLMs
Interpretation. Proposition 1 demonstrates that the type I and type II errors of our simple test decrease exponentially in the text length.
4.2 Statitical test for detection among multiple LLMs
Statistical test. Given a string $\mathbf{Y}_{N}$ with size $N$, we design a statistical test to detect whether the text is generated by one of the models $\mathcal{A}=\{A_{1},\ldots,A_{p}\}$ or one of models =$\mathcal{B}=\{B_{1},\ldots,B_{q}\}$ generated the text. The null hypothesis $\mathbf{H}_{0}$ is that the text $\mathbf{Y}_{N}$ is generated by one of the models in $\mathcal{B}$, and the alternative hypothesis $\mathbf{H}_{1}$ is that it is generated by one of the models in $\mathcal{A}$. We first calculate the random variables $Z^{M}_{n}=:-\log\big{(}p^{M}_{n}(Y_{n})\big{)}$, and sum $Z^{M}_{n}=:-\log\big{(}p^{M}_{n}(Y_{n})\big{)}$, for all models $M\in\mathcal{A}\cup\mathcal{B}$. Our test rejects the null hypothesis $\mathbf{H}_{0}$ if for some $A_{i}\in\mathcal{A}$, we have
$\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}<\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n};\qquad\forall B_{j}\in\mathcal{B}.$
Otherwise, our test accepts the null hypothesis.
Type I and type II errors. Type I error occurs when the test incorrectly concludes that the text is generated by one of the models in $\mathcal{A}$ when it is written by one of the models in $\mathcal{B}$, and Type II error happens when the test fails to identify that text is generated by one of the models in $\mathcal{A}$ and incorrectly concludes that it is generated by one of the models in $\mathcal{B}$. Similar to our test for the two model version, Assumption 2 must hold for us to quantify our model’s type I and type II errors.
Proposition 2.
If Assumptions 1 and 2 hold, the type I error for our statistical test is upper bounded by
$2|\mathcal{A}|\exp\bigg{[}-\frac{N(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}+2\exp\bigg{[}-\frac{N\epsilon_{1}/2}{c_{1}\log(K)}\min\bigg{(}1,\frac{\epsilon_{1}/2}{c_{1}\log(K)}\bigg{)}\bigg{]}.$
and the type II error for our statistical test is upper bounded by
with constants $c_{1}$, $c_{3}$, and $\epsilon$ as introduced in Theorems 1 and 2.
Proof.
See Section A.6. ∎
4.3 Statistical test for detection between an LLM and human
We need to make the following assumption to design our test.
Assumption 3.
(minimum tangible difference). We assume that if the generative and evaluator models are different, then
$\frac{1}{N}\sum_{n=3}^{N}\mathbb{E}\bigg{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{|}\textbf{Y}_{n-2}\bigg{]}\geq 4\log^{2}\big{(}K\big{)}.$
Assumption 3 ensures that the two models satisfy a minimum distance in terms of their expected KL divergence. Clearly, if models are the same (or very similar), then $A$ and $B$ impose the same (or almost the same) probability distributions $p_{N}^{A}(y_{n})$ over $\mathbf{Y}_{N}$, and hence KL divergence becomes zero. This makes differentiation impossible. Assumption 3 rules out cases where the two models impose very similar distributions over the text under evaluation.
Lemma 4.
Under assumption 3, for any positive constant $c_{4}\leq\log^{2}(K)/2$, we have
$\mathbb{P}\bigg{(}|h_{N}(B,A)(\mathbf{Y}_{N})-h_{N}(A,A)(\mathbf{Y}_{N})|\leq c_{4}\bigg{)}\leq 2\exp\bigg{[}-\frac{Nc_{4}}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{c_{4}}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$
Proof.
See Section A.7. ∎
Remark 2.
We note that the only application of Assumption 3 for establishing our results is that this assumption ensures that the statement $c_{4}\leq\log^{2}(K)/2$ in Lemma 4 holds. Our results remain true ( even without Assumption 3) if there exists $c_{4}>0$ such that if the generative and evaluator models are different, then
$\mathbb{P}\bigg{(}|h_{N}(B,A)(\mathbf{Y}_{N})-h_{N}(A,A)(\mathbf{Y}_{N})|\leq c_{4}\bigg{)}\leq 2\exp\bigg{[}-\frac{Nc_{4}}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{c_{4}}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$
Statistical test. Given a string $\mathbf{Y}_{N}$ with size $N$, for arbitrary constants $t<c_{4}\leq\log^{2}(K)/2$, we design a statistical test to detect whether the evaluator model $A$ generated the text. The null hypothesis $\mathbf{H}_{0}$ is that the text $\mathbf{Y}_{N}$ is not generated by the evaluator model $A$ (e.g., it is generated by another model $B$), and the alternative hypothesis $\mathbf{H}_{1}$ is that $\mathbf{Y}_{N}$ is generated by the evaluator model $A$. We first calculate the random variable $Z_{n}=:-\log\big{(}p^{A}_{n}(Y_{n})\big{)}$, and then we calculate the sum $\frac{1}{N}\sum_{n=1}^{N}Z_{n}$. Our test rejects the null hypothesis $\mathbf{H}_{0}$ in favor of the alternative $\mathbf{H}_{1}$ if
$\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}\leq t.$
Otherwise, our test accepts the null hypothesis.
Type I and type II errors. Type I error occurs when the test incorrectly concludes that the text is generated by the evaluator model $A$ when it is written by $B$, and Type II error happens when the test fails to identify that text is generated by the evaluator model $A$ and incorrectly concludes that it is not written by $A$.
Proposition 3.
For any $t\geq 0$, the type II error for our statitical test is upper bounded by
$2\exp\bigg{[}-\frac{Nt}{c_{1}\log(K)}\min\bigg{(}1,\frac{t}{c_{1}\log(K)}\bigg{)}\bigg{]},$
with $c_{1}$ as introduced in Theorem 1.
Additionally, if Assumptions 1 and 3 hold, then for any positive $t<c_{4}$, type I error for our statistical test is upper bounded by
$2\exp\bigg{[}-\frac{Nc_{4}}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{c_{4}}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}+2\exp\bigg{[}-\frac{N(c_{4}-t)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(c_{4}-t)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.,$
with $c_{3}$ as introduced in Theorem 2 and $c_{4}$ as introduced in Assumption 3.
Proof.
See Section A.8. ∎
Interpretation. Proposition 3 demonstrates that the type I and type II errors of our test for detecting LLM $A$ vs. not LLM $A$ (that includes human) decrease exponentially in the text length.

### v4

@@ 3 Concentration inequalities
In this section, we provide concentration inequalities to show that if the string is generated by a model $B$ and evaluated by a model $A$, then $\frac{1}{N}\sum_{n=1}^{N}Z_{n}^{A}$ converges to the average cross-entropy of the string under $B$ and $A$, $h_{N}(B,A)$, with a high probability. If the text is generated and evaluated by model $A$, then $\frac{1}{N}\sum_{n=1}^{N}Z_{n}^{A}$ converges to the average entropy of the string under $A$ with high probability. In Appendix A.1, we provide the proofs for the results in this section. These concentration bounds are the backbones of the statistical tests we design in section 4. For tractability, we make a parametric assumption on the probability laws $p^{A}_{k}(.)$ and $p^{B}_{k}(.)$.
Assumption 1.
We assume that there exists $\epsilon>0$ such that $p^{A}_{n}(y_{k}),p^{B}_{n}(y_{k})\notin(0,\epsilon)$.
Assumption 1 implies that models $A$ and $B$ either do not associate any probability to a token $y\in\mathcal{X}$, or they assign a probability of at least $\epsilon$. Our theoretical results depend only on $\log(\epsilon)$. Hence, our theoretical bounds primarily rely on a constant shift in the logarithmic scale. Note that Assumption 1 is not restrictive and is aligned with practice, as computers only allow for a limited range of representable numbers due to finite precision in floating-point arithmetic. Very small probabilities are either rounded to zero or set to a minimum threshold to maintain numerical stability in computations (18). The first outcome of assumption 1 is that
With $\mathbb{E}_{p_{n}^{B}}\big{[}Z_{n}^{A}\big{]}<\infty$, we define the zero-mean random variable $X_{n}^{A}=Z_{n}^{A}-\mathbb{E}_{p_{n}^{B}}[Z_{n}^{A}]$. Theorem 1 presents our concentration inequality, which we apply to design our statistical tests.
Theorem 1.
(a) If model $A$ is the same as model $B$, there exists a constant $c_{1}>0$ independent of the evaluator model $A$ such that for any $t>0$ we have
$\mathbb{P}\bigg{(}\frac{1}{N}\left|\sum_{n=1}^{N}X_{n}^{A}\right|\geq t\bigg{)}\leq 2\exp\left[-\frac{Nt}{c_{1}\log(K)}\min\left(1,\frac{t}{c_{1}\log(K)}\right)\right].$
(b) If model $A$ is not model $B$, there exists a constant $c_{3}>0$ independent of models $A$ and $B$ such that for any $t>0$ we have
[…]
@@
For proof of part (a), see Section A.1.3 and for proof of part (b), see Section A.1.4. ∎
Interpretation. Theorem 1 - (a) states that if the string is generated by the same model as the evaluator model $A$, then the log-perplexity of the string under $A$ converges to the average entropy of the string under $A$, except with an exponentially small probability in string length. Theorem 1 - (b) states that if model $B$ generates the text, then, except with an exponentially small probability in string length, the log-perplexity of the string under the evaluator model $A$ converges to the average cross-entropy of the string under $B$ and $A$.
Theorem 1 implies that for any evaluator model $A$, $\sum_{n=1}^{N}Z_{n}^{A}$ converges to $h_{N}(B,A)$. As such, for $A=B$, $\sum_{n=1}^{N}Z_{n}^{A}$ converges to $h_{N}(A,A)$. Hence, our main idea in designing a statistical test is to check if $\sum_{n=1}^{N}Z_{n}^{A}$ has converged to $h_{N}(A,A)$. Our next section formalizes this idea.
4 Statistical test
We design our statistical tests using the result in Theorem 1 and then evaluate type I (false positive) and type II (false negative) errors. In Section 4.1, we design a composite statistical test that determines whether a text is generated by a model that belongs to a set of models $\mathcal{A}$ or belongs to a disjoint set of models $\mathcal{B}$. In Section 4.2, we study the case where we do not have access to the log-probabilities for the generator model (e.g., text is written by an unknown LLM, a human, etc.) and design a composite test to identify whether the model $A$ generated the text or not. In Appendix A.2, we provide the proofs for the results in this section.
4.1 Attribution among multiple LLMs
Given a string $\mathbf{Y}_{N}$, we design a statistical test to detect whether one of the models in $\mathcal{A}=\{A_{1},\ldots,A_{p}\}$ (non-sanctioned) or one of the models in =$\mathcal{B}=\{B_{1},\ldots,B_{q}\}$ (in-house) generated the text. The null hypothesis $\mathbf{H}_{0}$ is that the text $\mathbf{Y}_{N}$ is generated by a model in $\mathcal{B}$. We first calculate the random variables $Z^{M}_{n}=:-\log\big{(}p^{M}_{n}(Y_{n})\big{)}$, and sum over $Z^{M}_{n}=:-\log\big{(}p^{M}_{n}(Y_{n})\big{)}$ for all models $M\in\mathcal{A}\cup\mathcal{B}$. Our test rejects the null $\mathbf{H}_{0}$ if for some $A_{i}\in\mathcal{A}$, we have
Type I error occurs when the test incorrectly concludes that the text is generated by one of the models in $\mathcal{A}$, and Type II error occurs when the test fails to identify that the text is generated by one of the models in $\mathcal{A}$. Our theoretical bounds require that the two models do not impose the same probability distribution on the string. Hence, we introduce Assumption 2 as a minimum distance between log-probabilities for the two models.
Assumption 2.
(minimum difference-MM). We assume that if the generative and evaluator models are different, for an arbitrarily small $\epsilon_{1}>0$, we have
$\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\geq\epsilon_{1}.$
Assumption 2 ensures that the two models satisfy a minimum distance in terms of their KL divergence over the generated text  [^2] Using KL divergence as a measure for comparing the learnt distribution from a text generation model to the distribution of another text generation model or the distribution of human-written text is common, e.g., see 48 and the references within.. Note that KL divergence, by definition, is a non-negative value that demonstrates the distance between the two distributions over the next word for the two models. Our results show that the type I and type II errors for our statistical test are approximately $\exp\big{(}O(-N\epsilon_{1})\big{)}$, which indicates that even for small values of $\epsilon_{1}$, which can converge to zero with the text length (for example, $\epsilon_{1}=O(N^{-1/2})$), our test provides exponentially small type I and type II errors in the text length.
Proposition 1.
If Assumptions 1 and 2 hold, the type I error for our test is upper bounded by
and the type II error for our test is upper bounded by
with constants $c_{1}$, $c_{3}$, and $\epsilon$ as introduced in Theorem 1.
@@ 4.1 Attribution among multiple LLMs
See Section A.2.1. ∎
4.2 Model $A$ or not model $A$?
Given a string $\mathbf{Y}_{N}$, for arbitrary constants $t$, we design a statistical test to detect whether the evaluator model $A$ generated the text. In hypothesis testing, the null hypothesis reflects the default/conservative assumption. So, if we accuse someone of having used LLM $A$, they are innocent until proven guilty. Accordingly, the null $\mathbf{H}_{0}$ is that the text $\mathbf{Y}_{N}$ is not generated by the evaluator model $A$; it is generated by any model $B$ that we do not have white-box access to, which could be a human or some other language generation process [^3] Recall from section 2.1 that an application of the Bayes’ rule yields that any text generation process corresponds to some sequential model $B$ regardless of whether tokens are sequentially generated. Our test does not require any information about $B$.. The alternative $A$ is that it is generated by $A$. We first calculate the random variable $Z_{n}^{A}=:-\log\big{(}p^{A}_{n}(Y_{n})\big{)}$, and then we calculate $\frac{1}{N}\sum_{n=1}^{N}Z_{n}^{A}$. Our test rejects the null if
[…]
Type I error occurs when the test mistakenly concludes that the text is generated by the evaluator model $A$, and Type II error happens when the test incorrectly concludes that the text is not written by $A$. To derive type I and type II error controls, we impose Assumption 3. Note that if the two distributions are identical, no test can distinguish them, and, therefore, some assumption about distributions is always needed for any test to work.
Assumption 3.
(minimum difference-HM). We assume that if the generative and evaluator models are different, then for some positive $\epsilon_{2}>0$
We make two points about this assumption: First, note that $h_{N}(A,A)$ is always upper bounded by $\log(|\mathcal{X}|)$. Second, an objective in designing LLMs is to reduce the expected log-perplexity, which is equivalent to the average entropy $h_{N}(A,A)$, e.g., for language models similar to GPT-3, $h_{N}(A,A)$ is less than 4 (47). Our theoretical results only require that $|h_{N}(B,A)-h_{N}(A,A)|>\epsilon_{2}$ for an arbitrary $\epsilon_{2}>0$, and Assumption 3 is a sufficient condition that ensures that an $\epsilon_{2}>0$ exists.
Proposition 2.
For any $t\geq 0$, the type II error for our test is upper bounded by
with $c_{1}$ as introduced in Theorem 1. Also, if Assumptions 1 and 3 hold, then for any positive numbers $t<\epsilon_{2}/2$, and for $c_{3}$ as introduced in Theorem 1, the type I error of our test is upper bounded by
@@ 4.2 Model $A$ or not model $A$?
See Section A.2.2. ∎
@@
Interpretation. Proposition 2 demonstrates that the type I and type II errors of our test decrease exponentially in the text length. We note that our theoretical results also apply to a one-sided test that rejects the null hypothesis when $\frac{1}{N}\sum_{n=1}^{N}Z_{n}^{A}-h_{N}(A,A)(\mathbf{Y}_{N})\leq t.$